OpenStudy (anonymous):

What is the value of g[f(4)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5?

OpenStudy (amistre64):

what does f(4) equal?

OpenStudy (anonymous):

find f(4)

OpenStudy (anonymous):

then substite it into g(x)

OpenStudy (anonymous):

I don't understand....

OpenStudy (amistre64):

when x=4, what does f(x) equal?

OpenStudy (amistre64):

f(x) = 2x + 1 when x=4 f(4) = 2(4) + 1 = ??

OpenStudy (anonymous):

9?

OpenStudy (amistre64):

well, if we follow the usual multiplictaion and addition rules, then yes. f(4) = 2(4)+1 = 9

OpenStudy (amistre64):

now, whereever you see an f(4), we can replace it by 9

OpenStudy (amistre64):

g(f(4)) = g(9) sooo, g(x) = 2x - 5, g(f(4)) = g(9) = 2(9)-5

OpenStudy (anonymous):

I don't understand the answer ._.

OpenStudy (amistre64):

you dont understand what 2(9) - 5 is???

OpenStudy (amistre64):

g(f(4)) is a notation that i think causes alot confusion. the use of xs in both cases used to throw me for a loop. essentially you have g(x) such that x = f(4), but f(4) = 9 so determine g(9). or, another method is to construct a new function by inserting f(x) int g(x) 2(f(x)) - 5 2(2x+1) - 5 4x+2- 5 4x - 3, is equivalent to g(f(x)); now when x=4 .... you get the same result but in a different construction: 4(4) - 3

OpenStudy (amistre64):

its simpler just to work them separately